Greg drove at a constant speed in a rainstorm for 282 miles. He took a​ break, and the rain stopped He then drove 204 miles at a speed that was 4 miles per hour faster than his previous speed. If he drove for 10 ​hours, find the​ car's speed for each part of the trip.
Answer by jorel1380(3719)

Let n be the speed Greg drove during the rainstorm. Then, his speed after the skies had cleared would be n+4. So:
282/n + 204/n+4=10
282n+1128+204n=10(nē+4n)
10nē-446n-1128=0
5nē-223n-564=0
(5n+12)(n-47)=0
n=-12/5, 47
Greg drove at 47 mph during the rain, and 51 mph after.
☺☺☺☺